# Exploring the difficulty of estimating win probability: a simulation study

### v4

@@ 1 Introduction
Win probability (${\mathsf{WP}}$) as a function of game-state is a canonical value function in sports analytics (Baumer et al.,, 2023). In-game win probability is the crux of strategic decision making––make the decision that maximizes win probability––and is central to live betting on game outcomes. Fourth-down decision making in American football is a prime example: choose between a conversion attempt, field goal attempt, and punt according to win probability (Brill et al.,, 2024).
@@
Notably, statistical ${\mathsf{WP}}$ models are widely used today by analysts of American football (Lock and Nettleton,, 2014; Baldwin,, 2021). For instance, they form the foundation of open-source fourth-down recommendations (Brill et al.,, 2024). They are widely assumed to be reasonable and trustworthy because of the modeling flexibility of machine learning algorithms. Hence, in this paper we focus on statistical win probability estimators. The binary win/loss outcome variable, however, is noisy and features a strong dependence structure. In particular, each play in the same game shares the same draw of the win/loss outcome. Accordingly, Brill et al., (2024) found that these estimators are subject to substantial uncertainty and produce wide confidence intervals, even when fit from a large dataset featuring $229,635$ first-down plays across $4,101$ games and $16$ years. We suspect statistical ${\mathsf{WP}}$ estimators have high variance, exacerbated by the dependence structure of observational play-by-play data.
@@ 2.1 Introducing random walk football
We begin by describing the rules of random walk football. Random walk football begins at midfield (yardline $L/2$, where $L$ is an even integer). Each play, the ball moves left or right by one yardline with equal probability. If the ball reaches the left end of the field (yardline 00), team one scores a touchdown, worth $+1$ point. If the ball reaches the right end of the field (yardline $L$), team two scores a touchdown, worth $-1$ point. The ball resets to midfield after each touchdown. After $T$ plays, the game ends. If the game is still tied after $T$ plays, a fair coin is flipped to determine the winner.
@@
$S_{g,t+1}:=\begin{cases}S_{gt}+1&\text{ if }X_{gt}+\xi_{gt}=0\text{ (TD)}\\ S_{gt}-1&\text{ if }X_{gt}+\xi_{gt}=L\text{ (opp. TD)}\\ S_{gt}&\text{ else.}\end{cases}$ (3)
@@
${\mathsf{WP}}(T+1,x,s)=\begin{cases}1&\text{if }s>0\\ 1/2&\text{if }s=0\\ 0&\text{if }s<0,\\ \end{cases}$ (6)
@@ 2.2 Random walk football play-by-play data
We want to assess the impact of the dependence structure of the response variable of observational football data on the accuracy of a statistical win probability estimator. To do so, we compare the accuracy of a ${\mathsf{WP}}$ estimator fit from datasets of varying degrees of dependence. We introduce a parameter $K$ that controls the degree of dependence of the win/loss outcome variable in a generated dataset: we keep a random subsample of $K$ plays per generated game. $K=1$––keep just $1$ randomly sampled play in each of the $G$ simulated games––reflects independence across all plays in the filtered dataset because each game is generated independently. When $K=1$, the outcome variable $y_{gt}$ reflects an independent draw of the win/loss outcome of the game since each play is filtered from a separate game. $K=T$––keep all $T$ plays in each of the $G$ simulated games––reflects full dependence within each game and is equivalent to the original dataset. When $K=T$, the outcome variable $y_{gt}$ for each play $t$ in game $g$ reflects the same draw of the win/loss outcome of the game. Integer values of $K$ between $1$ and $T$ reflect intermediate degrees of dependence because just $1<K<T$ plays per game share the same draw of the response variable. For concreteness, we visualize example datasets for $K=1$, $K=3$, and $K=T$ in Table 4.
@@ 2.3 Ways to estimate win probability
Finally, statistical models are fit entirely from historical data. Given the results of a set of observed plays, statistical models fit the relationship between a binary win/loss outcome variable and certain game-state variables using regression or machine learning approaches. Notably, these models are widely used today by analysts of American football (Lock and Nettleton,, 2014; Baldwin,, 2021). They form the foundation of open-source fourth-down recommendations (Brill et al.,, 2024). These models are popular due to the accessibility of rich publicly available data (e.g., play-by-play data from nflFastR Carl and Baldwin, (2022)) and off-the-shelf machine learning tools (e.g., ${\mathsf{XGBoost}}$ Chen and Guestrin, (2016)). They are widely assumed to be reasonable and trustworthy because of the modeling flexibility of machine learning algorithms. For these reasons, in this paper we focus on statistical win probability estimators. The binary win/loss outcome variable, however, is noisy and features a strong dependence structure–each play in the same game shares the same draw of the win/loss outcome. Accordingly, Brill et al., (2024) found that these estimators are subject to substantial uncertainty and produce wide confidence intervals, even when fit from a large dataset featuring $229,635$ first-down plays across $4,101$ games and $16$ years. We suspect statistical ${\mathsf{WP}}$ estimators have high variance, exacerbated by the dependence structure of observational play-by-play data.
@@ 2.4 Bias-variance decomposition
In Figure 2 we visualize the bias-variance decomposition of a win probability estimator fit from these three types of datasets as a function of $\zeta$. The $x$-axis is $\log_{4}(\zeta)$ because $1014^{6}=4,096\approx 4,$ 101 is the sample size of the historical dataset of American football plays. We see that it is much better to use all plays per game rather than one independent play per game. Despite the strong dependence structure, keeping all the plays elucidates information about the structure of the covariate space. We also see that it would be much better if the plays had independent outcomes. This suggests that the dependence structure reduces the effective sample size of the dataset. We explore the extent of this reduction in the next section.
@@
Figure 3: Bias ($y$-axis) versus time $n$ ($x$-axis) and score differential $s$ (color) at midfield (field position $x=2$) of win probability estimated from a $(G=4,101,\ T=56,\ K=T)$ dataset. The lines denote the average values across $M=100$ simulations and the shaded regions denote plus-minus twice the standard errors.
@@ 2.6 Coverage of bootstrapped confidence intervals
Each type of bootstrap produces $B$ bootstrapped datasets $\{{\mathscr{D}}_{\text{train}}^{(b)}\}_{b=1}^{B}$ from the training dataset ${\mathscr{D}}_{\text{train}}$. We then fit a win probability model to each bootstrapped dataset using ${\mathsf{XGBoost}}$, $\{\widehat{\mathsf{WP}}_{b}\}_{b=1}^{B}$. From these, we form a $90\%$ confidence interval for ${\mathsf{WP}}({\mathbf{x}})$ at game-state ${\mathbf{x}}$ by the $5^{th}$ and $95^{th}$ quantiles of $\{\widehat{\mathsf{WP}}_{b}({\mathbf{x}})\}_{b=1}^{B}$. Letting $B=101$ in this section, we form a $90\%$ confidence interval by $[\widehat{\mathsf{WP}}_{(6)}({\mathbf{x}}),\ \widehat{\mathsf{WP}}_{(96)}({\mathbf{x}})]$. To avoid substantially low coverage near the extremes (${\mathsf{WP}}({\mathbf{x}})\approx 0$ or ${\mathsf{WP}}({\mathbf{x}})\approx 1$), we widen our confidence intervals when $\widehat{{\mathsf{WP}}}({\mathbf{x}})<0.025$ to have a lower bound of 0 and when $\widehat{{\mathsf{WP}}}({\mathbf{x}})>0.975$ to have an upper bound of 1.
@@
standard bootstrap $0.60\pm 0.01$ $0.027\pm 0.0005$
cluster bootstrap $0.71\pm 0.01$ $0.036\pm 0.0004$
randomized cluster bootstrap $0.76\pm 0.01$ $0.042\pm 0.0003$
@@ 2.7 The fractional bootstrap
In practice, it is impossible to tune the bootstrap at all, using either the traditional method from DiCiccio and Efron, (1996) or the fractional bootstrap. This is because we need to know true win probability in order to tune the $\beta$ or $\phi$ values in these alternative bootstraps to achieve adequate coverage. While we know true win probability in our simulation setting, it is an unobservable quantity in real life. Coverage is not calculable for real data because the win/loss outcome is either 00 or $1$ and win probability estimates lie in $(0,1)$.

### v5

@@ 1 Introduction
Win probability ($\mathsf{WP}$}) as a function of game-state is a canonical value function in sports analytics (Baumer et al.,, 2023). In-game win probability is the crux of strategic decision making––make the decision that maximizes win probability––and is central to live betting on game outcomes. Fourth-down decision making in American football is a prime example: choose between a conversion attempt, field goal attempt, and punt according to win probability (Brill et al.,, 2025).
@@
Notably, statistical $\mathsf{WP}$} models are widely used today by analysts of American football (Lock and Nettleton,, 2014; Baldwin,, 2021). For instance, they form the foundation of open-source fourth-down recommendations (Brill et al.,, 2025). They are widely assumed to be reasonable and trustworthy because of the modeling flexibility of machine learning algorithms. Hence, in this paper we focus on statistical win probability estimators. The binary win/loss outcome variable, however, is noisy and features a strong dependence structure. In particular, each play in the same game shares the same draw of the win/loss outcome. Accordingly, Brill et al., (2025) found that these estimators are subject to substantial uncertainty and produce wide confidence intervals, even when fit from a large dataset featuring $229,635$ first-down plays across $4,101$ games and $16$ years. We suspect statistical $\mathsf{WP}$} estimators have high variance, exacerbated by the dependence structure of observational play-by-play data.
@@ 2.1 Introducing random walk football
We begin by describing the rules of random walk football. Random walk football begins at midfield (yardline $L/2$, where $L$ is an even integer). Each play, the ball moves left or right by one yardline with equal probability. If the ball reaches the left end of the field (yardline 0), team one scores a touchdown, worth +1+1 point. If the ball reaches the right end of the field (yardline $L$), team two scores a touchdown, worth 1-1 point. The ball resets to midfield after each touchdown. After $T$ plays, the game ends. If the game is still tied after $T$ plays, a fair coin is flipped to determine the winner.
@@
[…]
@@
[…]
@@ 2.2 Random walk football play-by-play data
We want to assess the impact of the dependence structure of the response variable of observational football data on the accuracy of a statistical win probability estimator. To do so, we compare the accuracy of a $\mathsf{WP}$} estimator fit from datasets of varying degrees of dependence. We introduce a parameter $K$ that controls the degree of dependence of the win/loss outcome variable in a generated dataset: we keep a random subsample of $K$ plays per generated game. K=1K=1––keep just $1$ randomly sampled play in each of the $G$ simulated games––reflects independence across all plays in the filtered dataset because each game is generated independently. When $K=1$, the outcome variable $y_{gt}$ reflects an independent draw of the win/loss outcome of the game since each play is filtered from a separate game. K=TK=T––keep all $T$ plays in each of the $G$ simulated games––reflects full dependence within each game and is equivalent to the original dataset. When $K=T$, the outcome variable $y_{gt}$ for each play $t$ in game $g$ reflects the same draw of the win/loss outcome of the game. Integer values of $K$ between $1$ and $T$ reflect intermediate degrees of dependence because just $1<K<T$ plays per game share the same draw of the response variable. For concreteness, we visualize example datasets for $K=1$, $K=3$, and $K=T$ in Table 4.
@@ 2.3 Ways to estimate win probability
Finally, statistical models are fit entirely from historical data. Given the results of a set of observed plays, statistical models fit the relationship between a binary win/loss outcome variable and certain game-state variables using regression or machine learning approaches. Notably, these models are widely used today by analysts of American football (Lock and Nettleton,, 2014; Baldwin,, 2021). They form the foundation of open-source fourth-down recommendations (Brill et al.,, 2025). These models are popular due to the accessibility of rich publicly available data (e.g., play-by-play data from nflFastR Carl and Baldwin, (2022)) and off-the-shelf machine learning tools (e.g., $\mathsf{XGBoost}$} Chen and Guestrin, (2016)). They are widely assumed to be reasonable and trustworthy because of the modeling flexibility of machine learning algorithms. For these reasons, in this paper we focus on statistical win probability estimators. The binary win/loss outcome variable, however, is noisy and features a strong dependence structure–each play in the same game shares the same draw of the win/loss outcome. Accordingly, Brill et al., (2025) found that these estimators are subject to substantial uncertainty and produce wide confidence intervals, even when fit from a large dataset featuring $229,635$ first-down plays across $4,101$ games and $16$ years. We suspect statistical $\mathsf{WP}$} estimators have high variance, exacerbated by the dependence structure of observational play-by-play data.
@@ 2.4 Bias-variance decomposition
In Figure 2 we visualize the bias-variance decomposition of a win probability estimator fit from these three types of datasets as a function of $\zeta$. The $x$-axis is $\log_{4}(\zeta)$ because $4^{6}=4,096\approx 4,101$ is the sample size of the historical dataset of American football plays. We see that it is much better to use all plays per game rather than one independent play per game. Despite the strong dependence structure, keeping all the plays elucidates information about the structure of the covariate space. We also see that it would be much better if the plays had independent outcomes. This suggests that the dependence structure reduces the effective sample size of the dataset. We explore the extent of this reduction in the next section.
@@
Figure 3: Bias ($y$-axis) versus time $n$ ($x$-axis) and score differential $s$ (color) at midfield (field position $x=2$) of win probability estimated from a (G=4,101,T=56,K=T)(G=4,101,\ T=56,\ K=T) dataset. The lines denote the average values across $M=100$ simulations and the shaded regions denote plus-minus twice the standard errors.
@@ 2.6 Coverage of bootstrapped confidence intervals
Each type of bootstrap produces $B$ bootstrapped datasets $\{{\mathscr{D}}_{\text{train}}^{(b)}\}_{b=1}^{B}$ from the training dataset ${\mathscr{D}}_{\text{train}}$. We then fit a win probability model to each bootstrapped dataset using $\mathsf{XGBoost}$}, $\{\widehat{\mathsf{WP}}_{b}\}_{b=1}^{B}$. From these, we form a $90\%$ confidence interval for ${\mathsf{WP}}({\mathbf{x}})$ at game-state $\mathbf{x}$} by the $5^{th}$ and $95^{th}$ quantiles of $\{\widehat{\mathsf{WP}}_{b}({\mathbf{x}})\}_{b=1}^{B}$. Letting $B=101$ in this section, we form a $90\%$ confidence interval by $\widehat{\mathsf{WP}}_{(6)}({\mathbf{x}}),\ \widehat{\mathsf{WP}}_{(96)}({\mathbf{x}})]$. To avoid substantially low coverage near the extremes (${\mathsf{WP}}({\mathbf{x}})\approx 0$ or ${\mathsf{WP}}({\mathbf{x}})\approx 1$), we widen our confidence intervals when $\widehat{{\mathsf{WP}}}({\mathbf{x}})<0.025$ to have a lower bound of 0 and when $\widehat{{\mathsf{WP}}}({\mathbf{x}})>0.975$ to have an upper bound of 1.
@@
standard bootstrap $60\pm 0.01 0.0$$027\pm 0.0005$
cluster bootstrap $71\pm 0.01 0.0$$036\pm 0.0004$
randomized cluster bootstrap $76\pm 0.01 0.0$$042\pm 0.0003$
@@ 2.7 The fractional bootstrap
In practice, it is impossible to tune the bootstrap at all, using either the traditional method from DiCiccio and Efron, (1996) or the fractional bootstrap. This is because we need to know true win probability in order to tune the $\beta$ or $\phi$ values in these alternative bootstraps to achieve adequate coverage. While we know true win probability in our simulation setting, it is an unobservable quantity in real life. Coverage is not calculable for real data because the win/loss outcome is either 0 or $1$ and win probability estimates lie in $(0,1)$.